Actions
Ops/Comm Scheduling meeting : 7/11/2017¶
Location : CR21 at 14:00 pm
Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517
- Actions:
- HN Reboot data mover
- Paolo add more disk space on MICEraid5
- Need to give Tony Millington heads up on survey requirements in Nov/Dec once the run plan has settled (3 Absorber changes)
- Ajit looking at producing automated cooling channel magnet ramp down procedure for use by MOM's
- DC Update
- Shutdown activities - the list so far
Name | Task | Time required | Done? |
Henry | Restart conventional magnet controls crate | 1-2 hours (?) | Done. |
Josef | Replace filters in 8 compressors on mezzanine | 1-2 days | Done. Issues with SSU C5. |
Ajit | Test AMI controllers | 0.5 - 1 day | Done. |
Mark | Partial warming of DS to clean vacuum space and recool | 10 days | Cooling |
Ajit/MOM | Review current alarms | 1 day | Went through alarms with CM, all understood, some state changes req'd |
Joe/Henry | Change target patch fibres | 0.5 days | In progress, Joe down Tues/Weds. Inspect in synch |
John/Colin | Hall AC work? | ? | Proceeding. |
Turn on tracker helium | Melissa and Durga | ||
Finalise tracker calibration | Melissa |
- Others?
Tracker - Check David has finished his work. Need calibration run next week. Melissa & Durga to help wrap this up.
- Next user cycle : 2017/03 : Nov 14 - Dec 20
- Shifts are now open.
Shift Block | # Open shifts / 6 | # data-taking hours / day |
1 | 2 | 16 |
2 | 2 | 16 |
3 | 3.4* | 8 |
4 | 0 | 24 |
5 | 1 | 16 |
6 | 2 | 16 |
7 | 0 | 24 |
*Have 10.4 open shifts / 42
*At the very least we need:
- 1 more shifter for 3 days in shift block 3 to match with Milorad
- 1 more shifter in shift block 5 to make it a 24 hour shift block
- Pulser run will be this week, nominally on Thursday 9th. Beam tuning and startup on Nov 14.
- Have we organised this with ISIS?
- AOB
Updated by Boyd, Steven over 5 years ago · 17 revisions